OpenStudy (angelberina):

TRUE OR FALSE: If f '(2) = 3.1 and g '(2) = 7.3, then the graph of f(x) + g(x) has slope 10.4 at x = 2.

OpenStudy (kevin):

h(x) = f(x) + g(x) h(2) = f(2) + g(2)

OpenStudy (danjs):

recall the limit properties? \[\lim_{x \rightarrow a}[f(x) \pm g(x)]=\lim_{x \rightarrow a}f(x) \pm \lim_{x \rightarrow a}g(x)\]

OpenStudy (danjs):

you can say the derivatives add from that (f + g) ' = f ' + g '

OpenStudy (kevin):

h'(2) = f'(2) + g'(2) substitute f'(2) with 3.1 and g'(2) with 7.3